# Temporal Feature Matters: A Framework for Diffusion Model Quantization

### v3

@@ G Deployment on Edge and Mobile Devices
More than focusing on very high-calculus devices in Sec. 5.3, we also deploy quantized models on the edge and mobile devices. Since our methods do not induce any operators that require specific hardware support, there is no constraint for deploying our models on devices that support normal uniform quantization. As illustrated in Tab V, we deploy Stable-Diffusion on Nvidia Jetson Orin Nano and iPhone 15 Pro Max. The quantized model exceptionally achieves 2.85×2.85$\times$2.85 × and 1.98×1.98$\times$1.98 × speedup, respectively.
TABLE V: Inference analysis of our quantized Stable-Diffusion employing Distrubance-aware Selection with 50 denoising time-steps and batch size of 1. We employ TensorRT [65] and Apple Core ML toolkit [66] for deployment on Jetson Orin Nano and iPhone 15 Pro Max, respectively. Each hardware has 8GB GPU memory. The speedup ratio depends on the utilized hardware and acceleration toolkits.
@@
[…]
@@ H Visualization Results
we present random samples derived from FP and w4a8 quantized diffusion models with a fixed random seed. These quantized models were created through our complete framework or previous methods. As shown from Fig. II to Fig. IX, our framework yields results that closely resemble those of the FP model, showcasing higher fidelity. Moreover, it excels in finer details, producing superior outcomes in some intricate aspects (zoom in to closely examine the relevant images).
@@
(b) Q-Diffusion (w4a8)
(c) PTQD (w4a8)
(d) Ours (w4a8)
Figure I: Random samples from w4a8 quantized and FP LDM-4 on CelebA-HQ $256\times 256$. The resolution of each sample is $256\times 256$.
@@
(f) Q-Diffusion (w4a8)
(g) PTQD (w4a8)
(h) Ours (w4a8)
Figure II: Random samples from w4a8 quantized and FP LDM-4 on FFHQ $256\times 256$. The resolution of each sample is $256\times 256$.
@@
(b) Q-Diffusion (w4a8)
(c) Ours (w4a8)
Figure III: Random samples from w4a8 quantized and FP DDIM on CIFAR-10 $32\times 32$. The resolution of each sample is $32\times 32$.
@@
(b) Q-Diffusion (w4a8)
(c) PTQD (w4a8)
(d) Ours (w4a8)
Figure IV: Random samples from w4a8 quantized and FP LDM-8 on LSUN-Churches $256\times 256$. The resolution of each sample is $256\times 256$.
@@
(f) PTQD (w4a8)
(g) Ours (w4a8)
Figure V: Random samples from w4a8 quantized and FP LDM-4 on LSUN-Bedrooms $256\times 256$. The resolution of each sample is $256\times 256$.
@@
(b) PTQD (w4a8)
(c) Ours (w4a8)
Figure VI: Random samples from w4a8 quantized and FP LDM-4 on ImageNet $256\times 256$. The resolution of each sample is $256\times 256$.
@@
“A large display of many different types of doughnuts. ” A thick crusted pizza just taken out of the oven. ” “A thick man in a white suit and tie wearing a name badge. ”
@@
Figure VII: Random samples from w4a8 quantized and FP Stable Diffusion-v1-4 on MS-COCO captions. The resolution of each sample is $512\times 512$. The images below the corresponding prompts are generated from FP (Left), Q-Diffusion (Middle), and our framework (Right).
@@
Figure VIII: Random samples from w4a8 quantized and FP SD-XL-turbo. The resolution of each sample is $512\times 512$. The images below the corresponding prompts are generated from FP (Left), Q-Diffusion (Middle), and our framework (Right).
@@
Figure IX: Random samples from w4a8 quantized and FP SD-XL. The resolution of each sample is $1024\times 1024$. The images below the corresponding prompts are generated from FP (Left), Q-Diffusion (Middle), and our framework (Right).

### v4

@@ G Deployment on Edge and Mobile Devices
TABLE V: Inference analysis of our quantized Stable Diffusion employing Distrubance-aware Selection with $50$ denoising timesteps and batch size of $1$. We employ TensorRT [64] and Apple Core ML toolkit [65] for deployment on Jetson Orin Nano and iPhone $15$ Pro Max, respectively. Each hardware has $8$GB of GPU memory. The speedup ratio depends on the utilized hardware and acceleration toolkits.
@@
8/8 821.15 9.93(1.98×)
More than focusing on very high-calculus devices in Sec. V-C, we also deploy quantized models on the edge and mobile devices. Since our methods do not induce any operators that require specific hardware support, there is no constraint for deploying our models on devices that support normal uniform quantization. As illustrated in Tab. V, we deploy Stable Diffusion on NVIDIA Jetson Orin Nano and iPhone $15$ Pro Max. The quantized model achieves $\mathbf{2.85\times}$ and $\mathbf{1.98\times}$ speedup, respectively.
@@ H Visualization Results
We present random samples derived from FP and W4444A8888 quantized diffusion models with a fixed random seed. These quantized models were created through our complete framework or previous methods. As shown from Fig. II to Fig. IX, our framework yields results that closely resemble those of the FP model, showcasing higher fidelity. Moreover, it excels in finer details, producing superior outcomes in some intricate aspects (zoom in to closely examine the relevant images).
@@
(b) Q-Diffusion (W4444A8888)
(c) PTQD (W4444A8888)
(d) Ours (W4444A8888)
Figure I: Random samples from W4444A8888 quantized and FP LDM-4 on CelebA-HQ $256\times 256$. The resolution of each sample is $256\times 256$.
@@
(f) Q-Diffusion (W4444A8888)
(g) PTQD (W4444A8888)
(h) Ours (W4444A8888)
Figure II: Random samples from W4444A8888 quantized and FP LDM-4 on FFHQ $256\times 256$. The resolution of each sample is $256\times 256$.
@@
(b) Q-Diffusion (W4444A8888)
(c) Ours (W4444A8888)
Figure III: Random samples from W4444A8888 quantized and FP DDIM on CIFAR-10 $32\times 32$. The resolution of each sample is $32\times 32$.
@@
(b) Q-Diffusion (W4444A8888)
(c) PTQD (W4444A8888)
(d) Ours (W4444A8888)
Figure IV: Random samples from W4444A8888 quantized and FP LDM-8 on LSUN-Churches $256\times 256$. The resolution of each sample is $256\times 256$.
@@
(f) PTQD (W4444A8888)
(g) Ours (W4444A8888)
Figure V: Random samples from W4444A8888 quantized and FP LDM-4 on LSUN-Bedrooms $256\times 256$. The resolution of each sample is $256\times 256$.
@@
(b) PTQD (W4444A8888)
(c) Ours (W4444A8888)
Figure VI: Random samples from W4444A8888 quantized and FP LDM-4 on ImageNet $256\times 256$. The resolution of each sample is $256\times 256$.
@@
“A large display of many different types of doughnuts. ” “A thick crusted pizza just taken out of the oven. ” “A thick man in a white suit and tie wearing a name badge. ”
@@
Figure VII: Random samples from W4444A8888 quantized and FP Stable Diffusion-v1-4 on MS-COCO captions. The resolution of each sample is $512\times 512$. The images below the corresponding prompts are generated from FP (Left), Q-Diffusion (Middle), and our framework (Right).
@@
Figure VIII: Random samples from W4444A8888 quantized and FP SD-XL-turbo. The resolution of each sample is $512\times 512$. The images below the corresponding prompts are generated from FP (Left), Q-Diffusion (Middle), and our framework (Right).
@@
Figure IX: Random samples from W4444A8888 quantized and FP SD-XL. The resolution of each sample is $1024\times 1024$. The images below the corresponding prompts are generated from FP (Left), Q-Diffusion (Middle), and our framework (Right).